OpenStudy (anonymous):

A new company will release 3 new products (A, B, C). 20% of buyers will buy A, 55% will buy B and 25% will buy C. People that buy A will buy 1.5 units, B 2.5 and C .75. What percent of total sales will be product B?

OpenStudy (anonymous):

ok so im guesssing it's a weighted average question: so this is how I THINK it works... Product A (1.5) x Product B (2.5)x Product C (.75)= 186.25 and so total of Product B/Total

OpenStudy (apoorvk):

okay so, suppose there are a hundred buyers then, 20 buy A, and each of them buys 1.5 units. So total units of A bought = 1.5 x 20 = 30 55 buy B, and each of the 55 buys 2.5 units of B. So total units of B bought = 55 x 2.5 =137.5 and, 25 buy C, each buying 0.75 units. So total units of C bought = 25 x 0.75 = 18.75 so total unit of A, B, C sold combined = 186.25 And yes, now you have figured it out. :D

OpenStudy (anonymous):

ok i just wanted to check my answer...it's great to have someone discuss it with. You're great, what grade are you in?

OpenStudy (apoorvk):

Right now am a high-school pass out, will join engineering school in some time. so no real greatness in solving those questions =]

OpenStudy (anonymous):

well goodluck, you definitely have a gift in looking at numbers and solving these type of problems, especailly because engineering doesn't focus too much on probabilities etc too do you want to do another question?

OpenStudy (apoorvk):

Well, yeah sure. (and am not really gifted lol, where I live there are loads of people with better talent than me :) )

OpenStudy (anonymous):

Well I hope you continue your studies in highschool so you can get into a good engineering college, you def. have a gift for numbers. Ok so the next one (i made this one up) It costs $1 for a plain pretzel (expense is .60, profit is .40), but customers want more salt which will cost an extra $.25 to make per pretzel. In the past, 100 pretzels were sold- calculate how many additional pretzels need to be sold to cover the costs of adding the extra expense of the salt.

OpenStudy (anonymous):

266.66 more pretzels is the answer i think 100 pretzels gives a profit of 40$ Therefore the formula is x= the number of pretzels sold x-(x)expense = profit 100-100(.60)= 40$ then to make 40$ profit, x-(x).85= 40 .15x=40 x=40/.15 and finally it would take 266.7 more pretzels to = 40$ revenue again

OpenStudy (anonymous):

am i correct???

OpenStudy (apoorvk):

Yes yes, @khizzzer you are correct. and a very apt solution.

OpenStudy (anonymous):

yea i hate it when i dont have a answer key

OpenStudy (anonymous):

calculate how many additional pretzels need to be sold to cover the costs of adding the extra expense of the salt. If the question asking for how many additional pretzels need to be sold the answer is not going to be 267-100??